92 81 65.5 Question 17 (1 point) The mean height of a basketball team is 6 feet with a standard deviation of 0.2 feet. The teams center is 6.9 feet tall. Find the centers z score. Is his score unusual? 3.83, no 4.95, yes 4.5, yes 4, no Question 18 (1 point) Saved If your score on your next statistics test is converted to a z score, which of these z scores would you prefer? 1.00
Answer #1

17

mean (mu) = 6
sd = 0.2
team center (x) = 6.9

z = (x-mu)/sd = (6.9-6)/0.2 = 4.5

we know, as an unusual score is either less than -2 or greater than 2 and this value is
greater than 2, so

option C) 4.5, yes